OpenStudy (dtan5457):

Silly question, but if x=-10 of the equation x^2, is it -100 or 100? Why is it sometimes 100 and sometimes -100...

OpenStudy (anonymous):

If the question is like this \[-(10)^2 = -100\] if it's \[(-10)^2 = 100\] because essentially \[(-10)^2 = (-10)(-10) \implies 100\]

OpenStudy (anonymous):

Hope that helped!

OpenStudy (dtan5457):

Oh, I see. Thank you!

OpenStudy (anonymous):

Np :)